# When “Competency” in Reasoning Opens the Door to Vulnerability: Jailbreaking LLMs via Novel Complex Ciphers

### v3

@@ Abstract
Recent advancements in Large Language Model (LLM) safety have primarily focused on mitigating attacks crafted in natural language or common ciphers (e.g. Base64), which are likely integrated into newer models’ safety training. However, we reveal a paradoxical vulnerability: as LLMs advance in reasoning, they inadvertently become more susceptible to novel jailbreaking attacks. Enhanced reasoning enables LLMs to interpret complex instructions and decode complex user-defined ciphers, creating an exploitable security gap. To study this vulnerability, we introduce Attacks using Custom Encryptions (ACE), a jailbreaking technique that encodes malicious queries with novel ciphers. Extending ACE, we introduce Layered Attacks using Custom Encryptions (LACE), which applies multi-layer ciphers to amplify attack complexity. Furthermore, we develop CipherBench, a benchmark designed to evaluate LLMs’ accuracy in decoding encrypted benign text. Our experiments reveal a critical trade-off: LLMs that are more capable of decoding ciphers are more vulnerable to these jailbreaking attacks, with success rates on GPT-4o escalating from 40% under ACE to 78% with LACE. These findings highlight a critical insight: as LLMs become more adept at deciphering complex user ciphers–many of which cannot be preemptively included in safety training–they become increasingly exploitable. [^1] Data and Source Code available at https://github.com/DivijH/jailbreak_cryptography
@@
When “Competency” in Reasoning Opens the Door to Vulnerability:
Jailbreaking LLMs via Novel Complex Ciphers
Divij Handa  Zehua Zhang  Amir Saeidi  Shrinidhi Kumbhar  Chitta Baral Arizona State University {dhanda, zzhan645, ssaeidi1, skumbha4, cbaral}@asu.edu
@@ 1 Introduction
Figure 1: Common ciphers such as Base64, which are widely used on the internet and can be easily incorporated into safety training, are decrypted by both older and newer LLMs. While older models, like GPT-4, struggle with deciphering novel user-defined ciphers, newer models like GPT-4o can decrypt such ciphers, making them vulnerable to potential jailbreaks.
Figure 2: Our first jailbreaking method, Attack using Custom Encryptions (ACE), uses novel user-designed ciphers to jailbreak LLMs, achieving a higher attack success rate (ASR) compared to both common and uncommon ciphers. Building upon ACE, Layered Attacks using Custom Encryptions (LACE), combines two encryption techniques, such as Word Substitution Cipher followed by Word Reversal Cipher. All ASRs displayed correspond to GPT-4o.
With the recent rise of Large Language Models (LLMs) (Xi et al., 2023; Yang et al., 2024) and their widespread availability to the public, there has been an increasing need to enhance the security and robustness of these models. To address these challenges, numerous strategies have been proposed, including red-teaming (Ganguli et al., 2022; Ge et al., 2023), reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022a, b), preference optimization (Rafailov et al., 2024; Saeidi et al., 2024; Liu et al., 2024), and perplexity-based defenses (Jain et al., 2023; Hu et al., 2023). These methods aim to align LLMs with ethical guidelines by refining their responses and improving safety.
@@
As the reasoning capabilities of LLMs increase, they become more capable of following complex user instructions and decoding ciphers that are not present in their training corpora. Prior works (Jiang et al., 2024; Wei et al., 2024; Inie et al., 2023) show that encoding queries using ciphers can jailbreak LLMs, but are restricted to widely-studied ciphers. In this paper, we build on these works, showing that novel user-created ciphers are not only understood by LLMs but can be used to jailbreak them. This directly challenges the early work done by Yuan et al. (2023), which suggested that such ciphers were ineffective for jailbreaking.
To systematically investigate this phenomenon, we first introduce CipherBench, a benchmark evaluating LLM’s ability to decode ciphers spanning three categories: common, uncommon, and novel user-created ciphers, based on how rarely they are discussed in cryptographic literature (Banoth and Regar, 2023; Biswas et al., 2019). We propose a jailbreak attack, Attack using Custom Encryptions (ACE), which encodes harmful queries in novel user-created ciphers. Our experiments reveal that both open-source and proprietary LLMs, including Llama-3.1-8B-Instruct (88% ASR), Llama-3.1-70B-Instruct (54% ASR) (Dubey et al., 2024), GPT-4o (40% ASR) (Achiam et al., 2023), and Gemini-1.5-Flash (66% ASR) (Reid et al., 2024) can be jailbroken using ACE.
Building upon ACE, we introduce a second jailbreak attack, Layered Attacks using Custom Encryptions (LACE), to examine the impact of sequentially layering multiple ciphers on LLM safety (Fig 2). This makes the encoding more complex to solve but our results show an increase in ASR for GPT-4o from 40% to 78%. This observation gives a key insight, “As LLMs become better at reasoning, they become more vulnerable to complex attacks”.
@@
    LACE Framework: We demonstrate that layering multiple ciphers (e.g. word-substitution + word-reversal) amplifies vulnerabilities, achieving a 78% ASR on GPT-4o – nearly doubling the effectiveness of single-cipher attacks.
@@
    The Security-Reasoning Tradeoff: LLMs which demonstrate a better performance on decoding ciphers, are more vulnerable to jailbreaking attacks using ACE and LACE. As progress in LLM reasoning continues and LLMs become better at decoding new ciphers, these ciphers can be used to jailbreak them.
@@ 3 CipherBench: Evaluating Decoding Capabilities of LLMs
In this section, we evaluate the ability of LLMs to decode sentences encrypted using ten different ciphers. We design a benchmark called CipherBench and assess the capabilities of four state-of-the-art LLMs. Details of the prompts for the encryption methods are provided in Appendix B.
@@ 3.1 Benchmark Design
We evaluate four LLMs–Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct, GPT-4o, and Gemini-1.5-Flash, on CipherBench to assess their performance in decoding the encrypted sentences. To quantify their effectiveness, we report the Decryption Success Rate (DSR), defined as the percentage of successfully decoded sentences:
@@ 3.2 Performance on CipherBench
Fig 3 and 4 summarizes the performance of all LLMs on CipherBench. Analyzing the results, we observe a significant decrease in DSR when the sentences are composed of random characters rather than standard English words, with an average drop of 35.23%. This finding suggests that the LLMs’ decoding capabilities rely heavily on the syntactic and semantic relationships between words, rather than on an inherent understanding of the underlying decryption algorithm. Furthermore, we notice that LLMs are 4.44% better at decoding questions compared to declarative statements, a trend consistent across all LLMs. Shorter queries are also 3.56% easier to decode than longer queries.
When examining the performance across different cipher categories, we find that the DSR for common encryption techniques is 29.4%, whereas uncommon techniques have a DSR of 14.62%, and novel, user-defined techniques have the lowest DSR of 14.06%. Specifically, among all ciphers, Base64 is the most accurately solved with a DSR of 38.65%, while the ASCII Art Cipher proves to be the most challenging, with a DSR of just 2.71%. These results support the hypothesis that LLMs perform better at decrypting encodings that are more prevalent in their pre-training data.
Figure 4: Decryption Success Rate (DSR) of LLMs across 10 ciphers present in CipherBench. Common ciphers have the highest DSR followed by uncommon ciphers followed by novel user-designed ciphers.
@@ 4 Experimental Setup for Jailbreaking
We use four distinct LLMs for our experiments, GPT-4o (Achiam et al., 2023), Gemini-1.5-Flash (Reid et al., 2024), Llama-3.1-8B-Instruct, and Llama-3.1-70B-Instruct (Dubey et al., 2024). These models were chosen because they have undergone comprehensive safety training, making them suitable for assessing the robustness of their safety measures against jailbreaking attempts. For the Gemini-1.5-Flash model, we configured the safety setting to “Block most”, the highest safety level available through the Gemini API. All the experiments for all LLMs are performed with greedy decoding (i.e. temperature 0).
@@
To evaluate our jailbreaking techniques, we utilize the AdvBench dataset developed by Zou et al. (2023). Specifically, and in line with prior studies (Zeng et al., 2024; Jiang et al., 2024; Chao et al., 2023; Wei et al., 2024), we extracted the same subset of 50 samples from AdvBench. These samples were manually categorized into 14 different unsafe categories, facilitating analysis of potential biases in the safety training of the LLMs. The categories are described in Appendix D. To further validate our results, we evaluate our jailbreaking methods on JailbreakBench (Chao et al., 2024). The results on JailbreakBench are discussed in Appendix A.
Additionally, we add priming sentences to all baselines and our methods. These sentences are designed to nudge LLMs toward producing an unsafe response. These priming sentences were generated using a few-shot prompting approach with GPT-4o. Detailed prompts are provided in Appendix C.
@@
[…]
@@
To determine whether an LLM’s response is harmful, we employed LLM-as-a-judge strategy. Specifically, we utilized GPT-4o-mini to classify responses as either “Safe” or “Unsafe”. The prompt and human correlation are presented in Appendix E. We report the Attack Success Rate (ASR), which quantifies the effectiveness of the jailbreaking technique and is calculated as:

### v4

# Reliable ML from Unreliable Data

@@ Abstract
Recent advancements in Large Language Model (LLM) safety have primarily focused on mitigating attacks crafted in natural language or common ciphers (e.g. Base64), which are likely integrated into newer models’ safety training. However, we reveal a paradoxical vulnerability: as LLMs advance in reasoning, they inadvertently become more susceptible to novel jailbreaking attacks. Enhanced reasoning enables LLMs to interpret complex instructions and decode complex user-defined ciphers, creating an exploitable security gap. To study this vulnerability, we introduce Attacks using Custom Encryptions (ACE), a jailbreaking technique that encodes malicious queries with novel ciphers. Extending ACE, we introduce Layered Attacks using Custom Encryptions (LACE), which applies multi-layer ciphers to amplify attack complexity. Furthermore, we develop CipherBench, a benchmark designed to evaluate LLMs’ accuracy in decoding encrypted benign text. Our experiments reveal a critical trade-off: LLMs that are more capable of decoding ciphers are more vulnerable to LACE, with success rates on gpt-oss-20b escalating from 60% under ACE to 72% with LACE. These findings highlight a critical insight: as LLMs become more adept at deciphering complex user ciphers–many of which cannot be preemptively included in safety training–they become increasingly exploitable. [^1] Data and Source Code available at https://github.com/DivijH/jailbreak_cryptography
@@ 1 Introduction
Figure 1: Our first jailbreaking method, Attack using Custom Encryptions (ACE), uses novel user-designed ciphers to jailbreak LLMs, achieving a higher attack success rate (ASR) compared to both common and uncommon ciphers. Building upon ACE, Layered Attacks using Custom Encryptions (LACE), combines two encryption techniques, such as Word Substitution Cipher followed by Word Reversal Cipher. All ASRs displayed correspond to gpt-oss-20b.
With the recent rise of Large Language Models (LLMs) (Xi et al., 2023; Yang et al., 2024) and their widespread availability to the public, there has been an increasing need to enhance the security and robustness of these models. To address these challenges, numerous strategies have been proposed, including red-teaming (Ganguli et al., 2022; Ge et al., 2023), reinforcement learning from human feedback (RLHF) (Christiano et al., 2017; Ouyang et al., 2022; Bai et al., 2022a, b), preference optimization (Rafailov et al., 2024; Saeidi et al., 2024; Liu et al., 2024b), and perplexity-based defenses (Jain et al., 2023; Hu et al., 2023). These methods aim to align LLMs with ethical guidelines by refining their responses and improving safety.
@@
As the reasoning capabilities of LLMs increase, especially with thinking LLMs (Jaech et al., 2024), they become more capable of following complex user instructions and decoding ciphers that are not present in their training corpora. Prior works (Jiang et al., 2024; Wei et al., 2024; Inie et al., 2023) show that encoding queries using ciphers can jailbreak LLMs, but are restricted to widely-studied ciphers. In this paper, we build on these works, showing that novel user-created ciphers are not only understood by LLMs but can be used to jailbreak them. This directly challenges the early work done by Yuan et al. (2023), which suggested that such ciphers were ineffective for jailbreaking.
To systematically investigate this phenomenon, we first introduce CipherBench, a benchmark evaluating LLM’s ability to decode ciphers spanning three categories: common, uncommon, and novel user-created ciphers, based on how rarely they are discussed in cryptographic literature (Banoth and Regar, 2023; Biswas et al., 2019). We propose a jailbreak attack, Attack using Custom Encryptions (ACE), which encodes harmful queries in novel user-created ciphers. Our experiments reveal that both open-source and proprietary LLMs, including gpt-oss-20b (60% ASR) (Agarwal et al., 2025), Llama-3.1-8B-Instruct (88% ASR), Llama-3.1-70B-Instruct (54% ASR) (Dubey et al., 2024), GPT-4o (40% ASR) (Achiam et al., 2023), and Gemini-1.5-Flash (66% ASR) (Reid et al., 2024) can be jailbroken using ACE.
Building upon ACE, we introduce a second jailbreak attack, Layered Attacks using Custom Encryptions (LACE), to examine the impact of sequentially layering multiple ciphers on LLM safety (Fig 1). This makes the encoding more complex to solve, but our results show an increase in ASR for most models, including GPT-5 (OpenAI, 2025), which goes from 0% with ACE to 8% with LACE. This observation gives a key insight, “As LLMs become better at reasoning, they become more vulnerable to more complex attacks”.
@@
    LACE Framework: We demonstrate that layering multiple ciphers (e.g., word-substitution + word-reversal) amplifies vulnerabilities, achieving an 8% ASR on GPT-5 and 72% ASR on gpt-oss-20b.
@@
    The Security-Reasoning Tradeoff: LLMs that demonstrate a better performance on decoding ciphers are more vulnerable to more complex jailbreaking attacks using LACE. As progress in LLM reasoning continues and LLMs become better at decoding new ciphers, such ciphers can be used in multiple layers to form more complex attacks.
@@ 3 CipherBench: Evaluating Decoding Capabilities of LLMs
In this section, we evaluate the ability of LLMs to decode sentences encrypted using ten different ciphers. We design a benchmark called CipherBench and assess the capabilities of five state-of-the-art LLMs. Details of the prompts for the encryption methods are provided in Appendix E.
@@ 3.1 Benchmark Design
We evaluate five LLMs–gpt-oss-20b, Llama-3.1-8B-Instruct, Llama-3.1-70B-Instruct, GPT-4o, and Gemini-1.5-Flash–on CipherBench to assess their performance in decoding the encrypted sentences  [^3] While we assume GPT-5 to perform the best on CipherBench, we omit running it on our benchmark due to limited budget.. All LLMs except gpt-oss-20b were run with a temperature of 0 to maintain reproducibility. We found gpt-oss-20b to be unstable at a temperature of 0, so we use a temperature of 0.8. To quantify their effectiveness, we report the Decryption Success Rate (DSR), defined as the percentage of successfully decoded sentences:
@@ 3.2 Performance on CipherBench
Fig 2 and 3 summarizes the performance of all LLMs on CipherBench. Analyzing the results, we observe a significant decrease in DSR when the sentences are composed of random characters rather than standard English words, with an average drop of 78.82%. This finding suggests that the LLMs’ decoding capabilities rely heavily on the syntactic and semantic relationships between words, rather than on an inherent understanding of the underlying decryption algorithm. Furthermore, we notice that LLMs are 4.44% better at decoding questions compared to declarative statements, a trend consistent across all LLMs. Shorter queries are also 18.63% easier to decode than longer queries.
When examining the performance across different cipher categories, we find that the DSR for common encryption techniques is 35.95%, whereas uncommon techniques have a DSR of 21.53%, and novel, user-defined techniques have a DSR of 21.87%. Specifically, among all ciphers, LeetSpeak is the most accurately solved with a DSR of 46.20%, while the ASCII Art Cipher proves to be the most challenging, with a DSR of just 3.80%. These results support the hypothesis that LLMs perform better at decrypting encodings that are more prevalent in their pre-training data.
Figure 3: Decryption Success Rate (DSR) of LLMs across 10 ciphers present in CipherBench
@@ 4 Experimental Setup for Jailbreaking
We use six distinct LLMs for our experiments: GPT-5 (OpenAI, 2025), gpt-oss-20b (Agarwal et al., 2025), GPT-4o (Achiam et al., 2023), Gemini-1.5-Flash (Reid et al., 2024), Llama-3.1-8B-Instruct, and Llama-3.1-70B-Instruct (Dubey et al., 2024). These models were chosen because they have undergone comprehensive safety training, making them suitable for assessing the robustness of their safety measures against jailbreaking attempts. For the Gemini-1.5-Flash model, we configured the safety setting to “Block most”, the highest safety level available through the Gemini API. Most LLMs are used with greedy decoding (i.e., temperature 0). GPT-5 is used with a temperature of 1  [^4] As of August 2025, GPT-5 only supports a temperature of 1 https://community.openai.com/t/gpt-5-models-temperature/1337957. gpt-oss-20b is unstable with a temperature of 0, so we use a temperature of 0.8.
@@
To evaluate our jailbreaking techniques, we utilize the AdvBench dataset developed by Zou et al. (2023). Specifically, and in line with prior studies (Zeng et al., 2024; Jiang et al., 2024; Chao et al., 2023; Wei et al., 2024), we extracted the same subset of 50 samples from AdvBench. These samples were manually categorized into 14 different unsafe categories, facilitating analysis of potential biases in the safety training of the LLMs. The categories are described in Appendix G. To further validate our results, we evaluate our jailbreaking methods on JailbreakBench (Chao et al., 2024). The results on JailbreakBench are discussed in Appendix D.
Additionally, we add priming sentences to all baselines and our methods. These sentences are designed to nudge LLMs toward producing an unsafe response. These priming sentences were generated using a few-shot prompting approach with GPT-4o. Detailed prompts are provided in Appendix F.
@@
    Upside Down Cipehr - In this uncommon cipher, each character in the sentence is turned upside down. For example, the word “Paper” becomes “ P a p e r ”. Although the transformed text remains human-readable and is not typically used in cryptography, the tokenization process in LLMs represents these characters very differently, making them effective as a cipher for LLMs.
@@
To determine whether an LLM’s response is harmful, as done by previous research (Liu et al., 2024a; Chao et al., 2024), we employed an LLM-as-a-judge strategy. Specifically, we utilized GPT-4o-mini to classify responses as either “Safe” or “Unsafe”. The prompt and human correlation are presented in Appendix H. We report the Attack Success Rate (ASR), which quantifies the effectiveness of the jailbreaking technique and is calculated as:
@@ 5 Results and Discussion
Attack Method GPT-4o GPT-5 gpt-oss-20b Llama-3.1-70B-Inst
DSR ASR DSR ASR DSR ASR DSR ASR
Priming Sentence - 0.00 - 0.00 - 0.04 - 0.48
PAIR - 0.02 - 0.00 - 0.00 - 0.16
AutoDAN - - - - - - - 0.84
ArtPrompt 0.00 0.16 0.06 0.04 0.06 0.54 0.00 0.14
LeetSpeak 0.92 0.24 0.64 0.00 0.78 0.30 0.72 0.44
Base64 1.00 0.10 0.66 0.00 0.56 0.36 0.78 0.44
ROT-13 0.98 0.24 0.80 0.00 0.62 0.48 0.00 0.28
Keyboard Cipher 0.06 0.16 0.66 0.00 0.70 0.54 0.00 0.06
Upside Down 0.72 0.30 0.72 0.00 0.58 0.60 0.00 0.06
Word Reversal 0.94 0.40 0.38 0.00 0.44 0.16 0.14 0.32
Grid Encoding 0.42 0.36 0.76 0.00 0.72 0.48 0.02 0.08
Word Substitution 0.98 0.40 0.58 0.00 1.00 0.26 0.82 0.54
Table 1: DSR and ASR for baseline methods and our ACE technique across various LLMs on AdvBench. A ‘-’ in the DSR column indicates that decoding the question was not required for that method. A ‘-’ in the ASR column signifies the technique applies only to open-source LLMs. The highest ASR for each model is highlighted in bold. The highest DSR is underlined. We omit AutoDAN for gpt-oss-20b due to a lack of support for this model.
Attack Method GPT-4o GPT-5 gpt-oss-20b Llama-3.1-70B
Word Substitution Cipher + DSR ASR DSR ASR DSR ASR DSR ASR
ROT-13 Sentence 0.80 0.72 0.72 0.00 0.74 0.56 0.00 0.44
Mappings 0.54 0.70 0.74 0.00 0.64 0.62 0.00 0.42
Both 0.50 0.60 0.72 0.06 0.66 0.60 0.00 0.04
Keyboard Sentence 0.14 0.42 0.78 0.04 0.80 0.62 0.00 0.34
Mappings 0.02 0.18 0.74 0.00 0.68 0.64 0.00 0.20
Both 0.00 0.08 0.82 0.04 0.86 0.72 0.00 0.22
Grid Encoding Sentence 0.26 0.30 0.76 0.00 0.78 0.70 0.00 0.40
Mappings 0.02 0.18 0.76 0.02 0.66 0.56 0.00 0.18
Both 0.04 0.12 0.78 0.08 0.60 0.66 0.00 0.10
Word Reversal Sentence 0.92 0.78 0.72 0.00 0.74 0.38 0.00 0.46
Mappings 0.80 0.64 0.72 0.04 0.54 0.58 0.04 0.44
Both 0.80 0.64 0.72 0.00 0.60 0.60 0.00 0.08
5.1 Effectiveness of ACE
Table 1 shows the DSR and ASR of our first jailbreaking method, ACE, along with seven baseline methods for 4 LLMs. The performance on the rest of the models is shown in Table 3. From the tables, we observe that common encryption techniques, previously explored by Yuan et al. (2023) and Wei et al. (2024), exhibit high decryption rates but result in low ASR. This confirms our hypothesis that existing encryption techniques can be easily incorporated into the safety training of LLMs, reducing their effectiveness as jailbreaking methods. In contrast, uncommon and novel techniques still result in an almost equal, if not better, attack success to common ciphers. Upon manually verifying some responses, we noticed that LLMs are smart enough to accurately decode parts of a sentence and still infer the meaning of the encoded query from a partially constructed sentence.
ACE Performance Across Categories
Fig. 4 shows a heatmap for GPT-4o, indicating the ASR for all encryptions introduced in our paper, dissected by the unsafe categories. We observe that the categories of Murder, Adult Content, and Financial are jailbroken with more success, indicating biases that could be present in the safety training of GPT-4o. In contrast, categories like Violence, Hate Speech, Self-Harm, and Stalking are more resistant to these attacks. The top four categories are 39.17% more prone to jailbreaking than the bottom four.
Performance of ACE scaled to Model Size and Inference Compute
Although GPT-5, GPT-4o and Gemini-1.5-Flash are commercial models and their parameter sizes are not publicly disclosed, it is widely hypothesized that their parameter counts are in the trillions. [^5] GPT-4o and Gemini 1.5 Pro: How the New AI Models Compare - CNET Since Gemini-1.5-Flash is smaller than Gemini-1.5-Pro, we also believe it to be smaller than GPT-4o but larger than Llama-3.1-70B-Instruct. From Table 3, we observe that as the LLM size increases, the DSR also increases. This higher rate of decryption leads to a higher ASR for many ciphers. This suggests that future models, likely to be more powerful and sophisticated, could be more susceptible to even more intricate ciphers, highlighting a parallel need for more sophisticated safety training to address these challenges.
Moreover, GPT-5 and gpt-oss-20b are thinking LLMs and use increased compute at inference time, representing a new scaling paradigm. We observe this paradigm to be more robust to jailbreaking attempts, with ACE being ineffective on GPT-5. gpt-oss-20b shows increased ASR, due to thinking of the model containing harmful advice, which is often marked as “Unsafe” by our evaluator.
5.2 Performance of LACE
Table 2 and 4 show the DSR and ASR of all the LLMs when attacked using LACE.
Change in DSR with Layering
Upon layering the encoding techniques, it becomes more complicated for the models to successfully decrypt the ciphers, resulting in a lower DSR—a trend consistent across all non-thinking LLMs. Thinking LLMs, on the other hand, result in a similar DSR, highlighting their advanced reasoning capabilities. Out of the four non-thinking LLMs used in our study, only GPT-4o shows consistent decryption performance, with the rest of the models falling behind significantly. GPT-4o, estimated to be the largest model among the four, highlights that bigger models are capable of solving even complex layered ciphers.
Encoding Sentence vs. Encoding Mapping
From Tables 2 and 4, we observe that when the prompt is presented with the harmful query encrypted with an additional encryption on top of the Word Substitution Cipher, it is easier to decode compared to when just the mappings of the Word Substitution Cipher are encrypted. We hypothesize that the model can correct itself by deducing the surrounding words when decoding the entire query, whereas it has to independently decode each word in the mapping when only the mapping is encrypted. When both the query and mappings are encoded, the model’s decoding capabilities drop significantly.
LACE improves ASR for better reasoners
Since GPT-4o, GPT-5, and gpt-oss-20b are the only models able to decode at a significant level successfully, we observe an increase in ASR for these LLMs. Specifically, for ROT-13 and Word Reversal Cipher, which are easier to solve, GPT-4o shows the highest ASR rate, up to 78%, along with 92% DSR. It almost doubles the ASR of 40% from using just one cipher. This finding demonstrates our argument that bigger and more sophisticated models are more likely to be jailbroken when decoding complex ciphers.
Smaller models, on the other hand, weren’t able to decrypt the sentences properly, but they still decoded enough words of the query to understand the meaning and trigger an attack. However, since their DSR is so low, the effectiveness of layering the ciphers is not observed on these models, suggesting the need to assign an appropriate level of cipher complexity based on the model’s decoding capabilities.
5.3 Over-defensiveness
Figure 5: Plot showing the number of instances that are over-defended across all five encryptions used in method ACE.
Fig 5 illustrates the encodings that are over-defensive on GPT-5, gpt-oss-20b, GPT-4o, and Gemini-1.5-Flash. The maximum over-defensiveness is shown by gpt-oss-20b with Grid Encoding, reaching 40% over-defended instances. Upon manually inspecting the responses, we find that the responses by gpt-oss-20b contain some text stating not to mention harmful instructions. Our evaluation done by GPT-4o-mini fails to notice this nuanced output carefully and labels this as overdefensive. For the rest of the ciphers, Gemini-1.5-Flash had a very low DSR, indicating poor decryption performance and resulting in a low over-defensive count. For GPT-4o, the Word Reversal Cipher and Word Substitution Cipher are the most successfully decoded ciphers and thus show the highest over-defensiveness. GPT-5 is the most robust LLM, returning the correct response for benign queries.
6 Conclusion
In this study, we demonstrated that as LLMs become more adept at reasoning, they become more vulnerable to more complex attacks. While traditional jailbreak attacks often rely on a single common cipher like Base64 or ROT-13, we build on these and introduce Attacks using Custom Encryptions (ACE) and Layered Attacks using Custom Encryptions (LACE). Within the ACE framework, we developed two novel ciphers—Grid Encoding and a Word-Substitution Cipher—that significantly increased the attack success rate (ASR) on GPT-4o (40%) and gpt-oss-20b (60%). Additionally, with LACE, we enhanced the ASR by layering the Word-Substitution Cipher with other methods such as Word Reversal, resulting in a 38% improvement on GPT-4o and 12% improvement on gpt-oss-20b. GPT-5, being the “smartest” LLM used in our study, only shows an ASR of 8% using LACE, and is the most robust LLM in this study. These findings highlight the urgent need for more robust security measures in LLMs to anticipate and defend against increasingly complex attack vectors.
Appendix A Ethics Statement
Appendix B Limitations & Future Work
While our methods are successful in jailbreaking LLMs and demonstrate the vulnerabilities that arise with the increasing reasoning capabilities of the model, it represents an early step in this area and has some limitations, including but not limited to the following:
  1. 1.
    Due to resource constraints, we restricted the exploration to three open-source LLMs and three proprietary LLMs. We would like to expand this study to more LLMs, with different architectures, to better assess the strengths and limitations of these models.
  2. 2.
  3. 3.
    In LACE, we restrict to a combination of two ciphers; in future work, we would like to extend this to multiple layers of ciphers to create more complex attacks.
  4. 4.
    We use GPT-4o-mini to evaluate whether the response is “Safe” or “Unsafe”. This makes our evaluation imperfect, as it misclassified some responses. More discussion is present in Appendix H.
  5. 5.
    Our work is limited to single-turn conversations, but in the future, we would like to expand to multi-turn conversations.
Appendix C Additional Results
Results on AdvBench for more LLMs: